A car accelerates from the rest to a speed of 10m/s in 4 seconds. It then decelerates to rest in 2 seconds.
A)draw the speed-time for the car?
B)work out the total distance travelled?
1
Expert's answer
2020-04-20T10:00:14-0400

Distance when car accelerates from the rest:

"s\begin{matrix} \\ 1\end{matrix}=\dfrac{a\times t^2}{2}=\dfrac{2.5\times 4^2}{2}=20(meters)"

Distance, when car accelerates to the rest:

"s\begin{matrix} \\ 2\end{matrix}=V\times t+\dfrac{a\times t^2}{2}=10\times 2-\dfrac{5\times2^2}{2}=10(meters)"

Together:

"s=s\begin{matrix} \\ 1\end{matrix}+s\begin{matrix} \\ 2\end{matrix}=30(metres)"